# Safe Learning for Contact-Rich Robot Tasks: A Survey from Classical Learning-Based Methods to Safe Foundation Models

### v1

@@ Abstract
Contact-rich tasks pose significant challenges for robotic systems due to inherent uncertainty, complex dynamics, and the high risk of damage during interaction. Recent advances in learning-based control have shown great potential in enabling robots to acquire and generalize complex manipulation skills in such environments, but ensuring safety, both during exploration and execution, remains a critical bottleneck for reliable real-world deployment.
This survey provides a comprehensive overview of safe learning-based methods for robot contact-rich tasks. We categorize existing approaches into two main domains: safe exploration, which focuses on minimizing the risk of unsafe actions during the learning phase, and safe execution, which ensures policy robustness and constraint satisfaction during deployment and interaction. We review key techniques, including constrained reinforcement learning, risk-sensitive optimization, uncertainty-aware modeling, control barrier functions, and model predictive safety shields, and highlight how these methods incorporate prior knowledge, task structure, and online adaptation to balance safety and efficiency.
A particular emphasis of this survey is on how these safe learning principles extend to and interact with emerging robotic foundation models, especially vision-language models (VLMs) and vision-language-action models (VLAs), which unify perception, language, and control for contact-rich manipulation. We discuss both the new safety opportunities enabled by VLM/VLA-based methods, such as language-level specification of constraints and multimodal grounding of safety signals, and the amplified risks and evaluation challenges they introduce. Finally, we outline current limitations and promising future directions toward deploying reliable, safety-aligned, and foundation-model-enabled robots in complex contact-rich environments. More details and materials are available at our Project GitHub Repository.
@@ 1 Introduction
Robots are increasingly expected to operate in unstructured and dynamic environments, where interaction with the physical world is inevitable [billard2019trends, ajoudani2018progress]. Among such interactions, contact-rich tasks, such as assembly, insertion, cutting, or tool use, are particularly challenging due to the complex, discontinuous dynamics and the need for precise force and motion control. Traditional model-based control methods often struggle in these scenarios due to inaccurate or hard-to-obtain models, while learning-based methods promise flexibility and adaptability by leveraging data-driven representations [elguea2023review].
However, applying learning-based methods to contact-rich tasks raises critical safety concerns [brunke2022safe]. Exploration during training may lead to unsafe actions that cause damage to the robot, the environment, or humans  [noseworthy2025forge]. Similarly, the execution of learned policies in a deployment must guarantee constraint satisfaction and robustness, especially under distributional shifts or unmodeled dynamics [zhang2024srl]. Therefore, safety must be addressed both during exploration and during execution to ensure the reliable operation of autonomous systems in the real world.
@@
This survey aims to provide a structured overview of the current landscape, foster a deeper understanding of safe learning in contact-rich robotics, and inspire new research toward safe and capable autonomous systems. In particular, it emphasizes the emerging role of VLMs and VLAs as robotic foundation models that unify perception, language, and control for contact-rich tasks. By framing safety challenges and design principles in the context of these models, the survey seeks to guide the development of VLM/VLA-based systems that are not only powerful and general, but also verifiably safe in real-world deployment.
@@
A number of recent surveys have explored topics at the intersection of safety, learning, and contact-rich robotic manipulation. However, none comprehensively address the unique challenges and opportunities posed by safe learning in contact-rich tasks. Our survey aims to bridge this gap by providing a focused, methodical overview of the state-of-the-art, specifically within this critical niche. Several surveys have investigated safe learning and RL more generally. Works such as [brunke2022safe, zhao2023state, gu2024review] provide comprehensive overviews of safety mechanisms in learning-based control and safe RL, including constraints, shielding, and exploration strategies. More specialized perspectives are offered in surveys on state-wise safety [zhao2023state], control barrier functions in RL [guerrier2024learning] and shielded RL [odriozola2023shielded]. While these surveys address safety in learning, they often neglect the complexities introduced by physical interaction and contact dynamics.
@@
From the perspective of robotic manipulation and contact, surveys such as [suomalainen2022survey, kroemer2021review] and  [xu2019compare] analyze manipulation strategies and control frameworks, including both model-based and model-free methods. Others review variable impedance control [abu2020variable] and compliant mechanisms [samadikhoshkho2024review], which are crucial for adapting to contact uncertainties but do not directly integrate learning-based safety frameworks. The application of RL in robotic assembly has also been reviewed in recent literature [stan2020reinforcement, elguea2023review, das2025towards] with emphasis on task success and sample efficiency. However, these surveys primarily focus on task performance and learning architectures, offering limited discussion on safety guarantees during exploration or deployment. Some surveys focused on contact-rich tasiks but lost safety highlight, such as  [Tsuji2025asu] only focus on imitation learning methods.
@@
More recent efforts begin to address physical risk and safety in the era of data-driven and foundation model-based robotics [kojima2025comprehensive, lamba2025alignment, xie2025towards, wabersich2023data], highlighting the growing relevance of scalable safety mechanisms. However, they often center on high-level policy alignment or general-purpose systems, lacking specificity in contact-rich manipulation contexts.
@@
Advanced agentic embodied AI sheds light on a generalist robotic system, while the safety issue remains due to a lack of physical interaction data. [salimpour2025towards] Safety for VLA robotic generalists lags because high-fidelity, failure and near‑miss contact data are scarce, and foundational semantic models still mis‑ground force, region, and tolerance constraints in physically complex interaction regimes [cruz2024learning]. Hybrid plan–parameterize–filter stacks that couple semantic/VLM layers with compliant action abstractions and certified online shields improve robustness under distribution shift but remain challenged by contact mode transitions and geometry/material variation [33, 330, 10705419, das2025towards]. Advancing VLA safety thus hinges on curated multimodal contact corpora, standardized force/violation metrics, and tighter analytical–data-driven integration for adaptive constraint grounding and reactive invariance maintenance [salimpour2025towards].
@@
Lastly, surveys on human-like manipulation [huang2025human] force/torque control [sy2025review], and safety in robotic manufacturing [lakshminarayanan2024robots] provide valuable domain-specific insights but remain largely disconnected from the safe learning perspective, especially the standard of safety was discussed in [lakshminarayanan2024robots].
@@
In contrast to the aforementioned surveys, our work uniquely focuses on the intersection of safe learning and contact-rich robotics, where safety must be ensured not only in policy learning but also in dynamic interaction with the environment. We systematically review methods that account for safety during physical contact, discuss open challenges in real-world applications, and outline promising directions for future research. We conducted a detailed analysis of the research intensity in this field over recent years, covering the period from 2018 to the present (2025). Fig. 1 shows the trend of research interest in this area.
@@
Figure 1: Trend of research publications with the keywords related to the safety, contact-rich robotics and learning. Representative example works are shown: contact-rich tasks, safe control, and learning-based methods [11023035, luu2025manifeel].
@@
Therefore, it is necessary to provide a comprehensive survey of recent advancements in safe learning for robotic contact-rich tasks. This survey focuses on learning-based approaches that directly integrate safety mechanisms into the learning process, such as safe reinforcement learning and model-based safety strategies. We deliberately exclude imitation learning methods, as their primary objective is to replicate demonstrated behaviors rather than actively managing safety during exploration or execution, an aspect critical in contact-rich environments where uncertainty and dynamic interaction play a significant role.
@@
    A Safety-Centric Taxonomy We introduce a structured taxonomy that categorizes safe learning approaches based on key dimensions, including learning phase (exploration vs. execution), level of safety integration (planning, control, or end-to-end), and modalities used (force/torque, vision, etc.). This survey provides a comprehensive lens through which researchers can analyze existing methods and identify safety design trade-offs.
@@
    Contextualization Within Contact-Rich Tasks Beyond general safe learning, we focus on its application to contact-rich robotic tasks such as insertion, polishing, and assembly. We detail how safety constraints are embedded in these tasks, and map the methods used to specific operational challenges (e.g., compliance, contact-inavatiable tasks, collision avoidance, and force control).
@@
    Identification of Gaps, Challenges, and Future Directions. We synthesize open research questions and outline critical challenges such as sim-to-real transfer under safety constraints, the scarcity of standardized benchmarks, and the need for provably safe generalization. We also discuss underexplored directions, including hybrid control-learning frameworks and human-in-the-loop safety mechanisms. Most importantly, we highlight the challenges and future opportunities in integrating safe contact-rich learning with large robotic foundation models, particularly VLMs and VLAs.
The remainder of this survey is organized as follows. In Section 2, we provide background on contact-rich tasks and formalize safety definitions in this context. Section 3 categorizes existing safe learning methods from multiple perspectives, including learning paradigms and safety mechanisms. In Section 4, we discuss open challenges and limitations in current approaches. Finally, Section 5 outlines promising future research directions to advance safe learning in robotic contact-rich tasks.
@@ 2 Background and Definitions in Safe Robotic Interaction
Figure 2: Illustrations of typical contact-rich tasks: Surface grinding [142], peg-in-hole [79], massaging [19] and non-prehensile object manipulation [kim_pre-_2023, lee_non-prehensile_2025]. The images are generated using ChatGPT 5.1.
@@ 2.1 Robotic Contact-Rich Tasks
The definition of contact-rich tasks varies in the literature, and it was also referred to by different terms such as physical interaction tasks [hogan2022contact], or manipulation in contact [suomalainen2022survey]. According to our definition, a task is contact-rich when its successful execution requires dynamic and extended physical contact with the environment.
In this regard, obstacle avoidance is not a contact-rich task: even though contact can happen sometimes, the task can be accomplished without contact. Simple pick-and-place tasks are not included even though they involve contacts with objects, because these contacts are fixed (i.e., not dynamic) after the grasp. However, tool-use tasks where the grasped object interacts with the environment, or in-hand manipulation where the contacts between the object and the robot hand are not fixed, are considered contact-rich. The definition also excludes the impact-based tasks such as throwing [100], hitting [48] or catching [zhang2025catch, 10611106] an object, in which the contact is momentary or sparse (i.e., not extended). These tasks are contact-intensive, but not contact-rich.
@@ 2.2 Definition of Safety in Contact-Rich Tasks
Safety in contact-rich manipulation spans multiple, complementary notions that arise from diverse objectives, stakeholders, and operational contexts [wabersich2023data]. Rather than a single definition, safety should be understood along several axes: physical interaction and energy exchange; stability and invariance; constraint satisfaction and reachability; risk- and data-driven guarantees; human-centered factors and standards [ISO15066]; and phase- and level-specific integration (exploration vs. execution; planning vs. control). We formalize these axes and situate them within contact-rich robotics with representative references.
@@ 2.2.1 Physical-Interaction Safety: Contact-force, Energy and Passivity
Contact force is a primary lens for safety in contact-rich manipulation. Excessive normal or tangential forces, whether as high transient impact at contact onset or as sustained forces/pressures during steady interaction, can damage workpieces and tooling, or injure humans in collaborative settings. A practical definition of safety, therefore, specifies force envelopes over time, including bounds on magnitude, rate of change, and impulse, together with region- or phase-specific limits (e.g., fragile surfaces, no-touch faces). Reviews on interaction/impedance control and force regulation formalize these aspects and their stability implications [ajoudani2018progress, calanca2015review].
@@
Achieving force-centric safety requires (i) reliable sensing/estimation of contact forces and states, (ii) controllers that prioritize force regulation under uncertainty, and (iii) runtime mechanisms that enforce hard limits. Force/torque sensing combined with observers and multimodal perception [huang_unified_2025] supports contact-state estimation (onset, sticking/slip) and force-grounded policy inputs; this enables early guard actions (e.g., approach deceleration, alignment) to reduce peaks and avoid lateral overload. Classical hybrid position–force control and impedance/admittance control provide physically interpretable means to track force references while maintaining compliant interaction; variable impedance schemes adapt stiffness/damping online to keep forces within bounds amid modeling error and task variation [ajoudani2018progress, calanca2015review]. In learning-integrated systems, producing high-level references (pose/force/impedance gains) rather than direct torques helps preserve stability margins while meeting force objectives [yuan2022safe].
@@
Enforcement at execution time benefits from casting force bounds (and friction cones) as constraints. Model-based safety filters, such as MPC with hard/soft constraints or barrier-based action projection, can minimally modify a nominal (possibly learned) command to preserve admissible contact forces, especially near transitions (free-space to contact, contact to sliding). This projection viewpoint also supports chance/risk-aware limits to reduce rare but damaging force spikes [yuan2022safe, 29, 152]. When specifications are given in natural language or task rules (e.g., “do not exceed 5 N on cable,” “avoid pressing the painted surface”), grounding them into formal force/region constraints and injecting them into the safety layer provides reliable guardrails during execution [33, 100].
@@
Energy and passivity complement the force-first view by regulating how much energy the robot injects into the environment. Passivity indices and energy reservoirs (energy tanks) offer monitors and buffers that cap impact energy and prevent force runaway under delays, unmodeled compliance, or high-gain settings. In practice, combining passivity-aware impedance/admittance with explicit force envelopes yields robust behavior across approach, impact, and sustained-contact phases, and remains compatible with learning-based references and uncertainty-aware shields [ajoudani2018progress, calanca2015review, yuan2022safe].

### v2

@@ Abstract
Contact-rich tasks pose significant challenges for robotic systems due to inherent uncertainty, complex dynamics, and the high risk of damage during interaction. As learning-based methods increasingly empower robots to acquire and generalize sophisticated manipulation skills in these challenging settings, the need for principled and comprehensive safe learning has become more urgent. Motivated by this need, we present the first safety-centric survey that places safety at the core of learning for contact-rich robotic tasks, examining how safety principles are formulated, enforced, and evaluated across classical learning-based approaches and emerging robotic foundation models. We categorize existing approaches into two main domains: safe exploration, which focuses on minimizing the risk of unsafe actions during the learning phase, and safe execution, which ensures policy robustness and constraint satisfaction during deployment and interaction. We review key techniques, and highlight how these methods incorporate prior knowledge, task structure, and online adaptation to balance safety and efficiency. A particular emphasis of this survey is on how these safe learning principles extend to and interact with emerging robotic foundation models, especially vision-language models (VLMs) and vision-language-action models (VLAs). We discuss both the new safety opportunities enabled by VLM/VLA-based methods, such as language-level specification of constraints and multimodal grounding of safety signals, and the amplified risks and evaluation challenges they introduce. Finally, we outline current limitations and promising future directions toward deploying reliable, safety-aligned, and foundation-model-enabled robots in complex contact-rich environments. More details and materials are available at our Project GitHub Repository.
@@ 1 Introduction
Robots are increasingly expected to operate in unstructured and dynamic environments, where interaction with the physical world is inevitable [29, 6]. Among such interactions, contact-rich tasks such as assembly, insertion, cutting, or tool use are particularly challenging due to their complex, discontinuous dynamics and the need for precise force and motion control. Traditional model-based control methods often struggle in these scenarios due to inaccurate or hard-to-obtain models, while learning-based methods promise flexibility and adaptability by leveraging data-driven representations [70].
However, applying learning-based methods to contact-rich tasks raises critical safety concerns. Exploration during training may lead to unsafe actions that damage the robot, the environment, or nearby humans. Similarly, the execution of learned policies in deployment must guarantee constraint satisfaction and robustness, especially under distributional shifts. Therefore, safety must be addressed both during exploration and during execution to ensure the reliable operation of autonomous systems in the real world.
@@
This survey aims to provide a structured overview of the current landscape, foster a deeper understanding of safe learning in contact-rich robotics, and inspire new research toward safe and capable autonomous systems. In particular, it emphasizes the emerging role of vision-language models (VLMs) and vision-language-action models (VLAs) as robotic foundation models that unify perception, language, vision, and control, but safey in contact-rich tasks is still an open challenge [307, 371]. By framing safety challenges and design principles in the context of these models, the survey seeks to timely guide the development of VLM/VLA-based systems that are not only powerful and general but also verifiably safe in real-world deployment.
1.1 Related Surveys and Their Limitations
@@
A number of recent surveys have explored topics at the intersection of safety, learning, and contact-rich robotic manipulation. However, none comprehensively address the unique challenges and opportunities posed by safe learning in contact-rich tasks. Our survey aims to bridge this gap by providing a focused, methodical overview of the state-of-the-art, specifically within this critical niche.
Several surveys have investigated safe learning and reinforcement learning (RL) more generally. Works such as [39, 383, 92] provide comprehensive overviews of safety mechanisms in learning-based control and safe RL, including constraints, shielding, and exploration strategies. More specialized perspectives are offered in surveys on state-wise safety [383], control barrier functions in RL [93] and shielded RL [229]. While these surveys address safety in learning, they often neglect the complexities introduced by physical interaction and contact dynamics.
@@
From the perspective of robotic manipulation and contact, surveys such as [291, 144] and  [344] analyze manipulation strategies and control frameworks, including both model-based and model-free methods. Others review variable impedance control [2] and compliant mechanisms [254], which are crucial for adapting to contact uncertainties but do not directly integrate learning-based safety frameworks.
The application of RL in robotic assembly has also been reviewed in recent literature [285, 70, 62] with emphasis on task success and sample efficiency. However, these surveys primarily focus on task performance and learning architectures, offering limited discussion on safety guarantees during exploration or deployment. A recent survey paper [310] reviews the imitation learning methods for contact-rich robotic tasks. Differently, we center our review specifically on online safe learning methods, without focusing on imitation learning.
@@
More recent efforts begin to address physical risk and safety in the era of data-driven and foundation model-based robotics [140, 150, 340, 314], highlighting the growing relevance of scalable safety mechanisms. However, they often center on high-level policy alignment or general-purpose systems, lacking specificity in contact-rich manipulation contexts.
@@
This survey also concentrates on the key safety gaps and data requirements underlying robot foundation models [307, 367]. Advanced agentic embodied AI sheds light on a generalist robotic system, while the safety issue remains due to a lack of physical interaction data. [252] Safety for VLA robotic generalists lags because high-fidelity, failure and near‑miss contact data are scarce, and foundational semantic models still mis‑ground force, region, and tolerance constraints in physically complex interaction regimes [57]. Hybrid plan–parameterize–filter stacks that couple semantic/VLM layers with compliant action abstractions and certified online shields improve robustness under distribution shift but remain challenged by contact mode transitions and geometry/material variation [367, 307, 388, 62]. Advancing VLA safety thus hinges on curated multimodal contact corpora, standardized force/violation metrics, and tighter analytical–data-driven integration for adaptive constraint grounding and reactive invariance maintenance [252].
@@
Lastly, surveys on human-like manipulation [115] force/torque control [293], and safety in robotic manufacturing [149] provide valuable domain-specific insights but remain largely disconnected from the safe learning perspective, especially the standard of safety was discussed in [149].
1.2 Our Scope and Contribution
@@
In contrast to the aforementioned surveys, our work uniquely focuses on the intersection of safe learning and contact-rich robotics, where safety must be ensured not only in policy learning but also in dynamic interaction with the environment. We systematically review methods that account for safety during physical contact, discuss open challenges in real-world applications, and outline promising directions for future research. We conducted a detailed analysis of the research intensity in this field over recent years, covering the period from 2018 to the present 2025. Figure 1 illustrates the trend of research interest in this area, along with the examples of robotic experiments associated with the keywords “safety,” “contact-rich robotics,” and “learning.”
@@
Figure 1: Trend of research publications that include keywords related to both safety, contact-rich robotics and learning. Representative example works are shown: contact-rich tasks, safe control, and learning-based methods [264, 198].
@@
Therefore, it is necessary to provide a comprehensive and timely survey of recent advancements in safe learning for robotic contact-rich tasks. This survey focuses on learning-based approaches that directly integrate safety mechanisms into the learning process, such as safe reinforcement learning and model-based safety strategies. We deliberately exclude imitation learning methods, as their primary objective is to replicate demonstrated behaviors rather than actively managing safety during exploration or execution, an aspect critical in contact-rich environments where uncertainty and dynamic interaction play a significant role. Regarding the scope of task, see 2.1.
@@
    A safety-centric taxonomy. We introduce a structured taxonomy that categorizes safe learning approaches by key dimensions, including learning phase (exploration vs. execution), level of safety integration (planning, control, or end-to-end), and modalities used (force/torque, vision, etc.). This survey provides a comprehensive lens for analyzing existing methods and identifying safety design trade-offs.
@@
    Contextualization within contact-rich tasks. Beyond general safe learning, we focus on its application to contact-rich robotic tasks such as insertion, polishing, and assembly. We detail how safety constraints are embedded in these tasks, and map the methods used to specific operational challenges (e.g., compliance, contact-inavatiable tasks, collision avoidance, and force control).
@@
    Identification of gaps, challenges, and future directions. We synthesize open research questions and outline critical challenges such as sim-to-real transfer under safety constraints, the scarcity of standardized benchmarks, and the need for provably safe generalization. We also discuss underexplored directions, including hybrid control-learning frameworks and human-in-the-loop safety mechanisms. Most importantly, we highlight the challenges and future opportunities in integrating safe contact-rich learning with large robotic foundation models, particularly VLM and VLA.
1.3 Review Methodology
To ensure a comprehensive and systematic review of safe learning methods for robotic contact-rich tasks, we employed a multi-step methodology:
  • Literature investigation: We conducted an extensive investigation across major academic databases, including IEEE Xplore, ACM Digital Library, SpringerLink, and Google Scholar. Keywords used in the literature search included combinations of ”safe”, ”contact-rich tasks” and ”learning”. To capture the most recent advancements, specifically, we use the following compliant search string: (”robot” OR ”robotic”) AND (”safe” OR ”safety” OR ”stable”) AND (”contact-rich” OR ”multi-contact” OR ”contact-intensive” OR ”physical interaction” OR ”environment interaction” OR ”contact interaction” OR ”contact manipulation” OR ”contact force”) AND (intitle:”learning” OR intitle:”learn”).
  • Inclusion criteria: We focused on both peer-reviewed and pre-print articles published from 2018 to 2025 that specifically addressed and/or discussed safety in learning-based approaches for contact-rich robotic tasks. Both theoretical contributions and empirical studies with experimental validation were considered.
  • Analysis: Each paper was analyzed for its methodological contributions, safety guarantees and discussion among multiple different perspectives, experimental setups including simulation and real-world experiments, and limitations.
The remainder of this survey is organized as follows. In Section 2, we provide background on contact-rich tasks and formalize safety definitions in this context. Section 3 categorizes existing safe learning methods from multiple perspectives, including learning paradigms and safety mechanisms. In Section 4, we discuss open challenges and limitations in current approaches. Finally, Section 5 outlines promising future research directions to advance safe learning in robotic contact-rich tasks. Especially, we highlight the challenges and perspectives in large robot foundation models, such as VLM/VLA.
@@ 2 Background and Definitions in Safe Robotic Interaction
Figure 2: Illustrations of typical contact-rich tasks: Surface grinding, peg-in-hole, massaging, and non-prehensile object manipulation. The images are generated using ChatGPT 5.1.
@@ 2.1 Robotic Contact-Rich Tasks
The definition of contact-rich tasks varies in the literature, and it was also referred to by different terms such as physical interaction tasks [104], or manipulation in contact [291]. According to our definition, a task is contact-rich when its successful execution requires dynamic and sustained physical contact with the environment, where motion and force are tightly coupled through contact constraints.
In this regard, for instance, obstacle avoidance is not a contact-rich task: even though contact can happen sometimes, the task can be accomplished without contact. Simple pick-and-place tasks are not included even though they involve contacts with objects, because these contacts are fixed (i.e., not dynamic) after the grasp. However, tool-use tasks where the grasped object interacts with the environment, or in-hand manipulation where the contacts between the object and the robot hand are not fixed, are considered contact-rich. The definition also excludes the impact-based tasks such as throwing [14], hitting [54] or catching [297] an object, in which the contact is momentary or sparse (i.e., not extended). These tasks are dynamic, but not contact-rich.
@@ 2.2 Definition of Safety in Contact-Rich Tasks
Safety in contact-rich manipulation spans multiple, complementary notions that arise from diverse objectives, stakeholders, and operational contexts [314]. Rather than a single definition, safety should be understood along several axes: physical interaction and energy exchange; stability and invariance; constraint satisfaction and reachability; risk- and data-driven guarantees; human-centered factors and standards [119]; and phase- and level-specific integration (exploration vs. execution; planning vs. control). We formalize these axes and situate them within contact-rich robotics with representative references.
@@ 2.2.1 Physical-Interaction Safety: Contact-force, Energy and Passivity
Contact force is a primary lens for safety in contact-rich manipulation. Excessive normal or tangential forces, whether as high transient impact at contact onset or as sustained forces/pressures during steady interaction, can damage workpieces and tooling, or injure humans in collaborative settings. A practical definition of safety, therefore, specifies force envelopes over time, including bounds on magnitude, rate of change, and impulse, together with region- or phase-specific limits (e.g., fragile surfaces, no-touch faces). Reviews on interaction/impedance control and force regulation formalize these aspects and their stability implications [6, 41].
@@
Achieving force-centric safety requires (i) reliable sensing/estimation of contact forces and states, (ii) controllers that prioritize force regulation under uncertainty, and (iii) runtime mechanisms that enforce hard limits. Force/torque sensing combined with observers and multimodal perception [116] supports contact-state estimation (onset, sticking/slip) and force-grounded policy inputs; this enables early guard actions (e.g., approach deceleration, alignment) to reduce peaks and avoid lateral overload. Classical hybrid position–force control and impedance/admittance control provide physically interpretable means to track force references while maintaining compliant interaction; variable impedance schemes adapt stiffness/damping online to keep forces within bounds amid modeling error and task variation [6, 41]. In learning-integrated systems, producing high-level references (pose/force/impedance gains) rather than direct torques helps preserve stability margins while meeting force objectives [6, 362].
@@
Enforcement at execution time benefits from casting force bounds (and friction cones) as constraints. Model-based safety filters, such as MPC with hard/soft constraints or barrier-based action projection, can minimally modify a nominal (possibly learned) command to preserve admissible contact forces, especially near transitions (free-space to contact, contact to sliding). This projection viewpoint also supports chance/risk-aware limits to reduce rare but damaging force spikes [362, 132, 123]. When specifications are given in natural language or task rules (e.g., “do not exceed 5 N on cable,” “avoid pressing the painted surface”), grounding them into formal force/region constraints and injecting them into the safety layer provides reliable guardrails during execution [367, 14].
@@
Energy and passivity complement the force-first view by regulating how much energy the robot injects into the environment. Passivity indices and energy reservoirs (energy tanks) offer monitors and buffers that cap impact energy and prevent force runaway under delays, unmodeled compliance, or high-gain settings. In practice, combining passivity-aware impedance/admittance with explicit force envelopes yields robust behavior across approach, impact, and sustained-contact phases, and remains compatible with learning-based references and uncertainty-aware shields [6, 41, 362].